OpenStudy (anonymous):

evaluate r+s/6 for r=15 and s=15

OpenStudy (anonymous):

15+15/6 = 15+2.5=17.5

OpenStudy (anonymous):

oh so simple, do solve it urself

OpenStudy (anonymous):

would the answer be 35/2

OpenStudy (anonymous):

i guess 6 is under r+s?

OpenStudy (anonymous):

or just under s?

OpenStudy (anonymous):

yes 35/2=17.5

OpenStudy (anonymous):

i said if the sume of the digits of a number is divisible by 3 then that number is divisible by 3 like 105 1+0+5 =6 6 is divisible by 3 so 105 is divisible by 3 105/6 and simplify = 35/2 correct

OpenStudy (anonymous):

mrjoselson which is it 35/2 or 17.5

OpenStudy (anonymous):

yes it is correct and it can be expressed as 17,5 in decimal form

OpenStudy (anonymous):

so Harkirat if the problem states simpify only should I write 35/2 or 17.5

OpenStudy (anonymous):

both r acceptable

OpenStudy (anonymous):

if it is from chapter of fractions, then 35/2 or 17 1/2 will do if it is not specifically from the chapter on fractions then any answer will do...